Sol Summary – April 18th

Crew 192 Sol Summary Report 18Apr2018

Sol 10

Summary Title: Canyons, Baby

Author: Victoria Varone

Mission Status: A beautiful day for the crew, mission nearing its end.

Sol Activity Summary: The crew went on two amazing EVAs today, exploring some of the major Martian canyons in the area. Lith Canyon and Murphy Canyon provided some of the most breathtaking views we’ve seen so far.

Then we get a little sun-gazing in at the Musk Observatory before cooking an awesome dinner for the crew that was a wonderful example of a team effort.

Look Ahead Plan: Aiming to make our last few EVAs fantastic ones and get in some great stargazing before we leave.

Anomalies in Work: None

Weather: Freezing in the morning, a bit warmer throughout midday and the afternoon. Calm with little wind and partly cloudy.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report, Astronomy Report

Support Requested: Please add Doug Campbell, incoming with Crew 193, to the Mission Support list

Maintenance request: See Operations Report

Sol Summary – April 17th

Crew 192 Sol Summary Report 17Apr2018

Sol 9

Summary Title: Don’t Feed the Penguin

Author: Victoria Varone

Mission Status: Nearing its end, aiming to fit in as much exploration as possible.

Sol Activity Summary: The crew went on two great EVAs today, exploring parts of the Martian terrain that we haven’t yet seen. It’s fascinating to see just how many different kinds of soil, rock, and sediment you can find even in just one area, in varying colors, textures, shapes, and sizes. The winds were strong today (and last night) so extra caution was exercised during EVAs.

Look Ahead Plan: Trying to fit in as many EVAs as possible (pending weather) since weekend weather may not be suitable and our mission ends on Saturday.

Anomalies in Work: None

Weather: Chilly, mix of clouds and sun, strong winds throughout the night and most of the day.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report

Support Requested: None

Maintenance request: See Operations Report

Sol Summary – April 16th

Crew 192 Sol Summary Report 16Apr2018

Sol 8

Summary Title: Houston, We Have a [Mouse] Problem

Author: Victoria Varone

Mission Status:

Sol Activity Summary: Our day started with a visit from a tiny companion ‒ code name “Paul Watney” in honor of our missing crew member and “The Martian” ‒ who made his way into the upper deck of the hab. He accompanied us in the hab for the day but then decided to return to Mars. (see Photo Report)

Our first EVA of the day was a long walking hike along the ridge. We climbed some of the dunes and hills and had to get creative along the way as there weren’t outlets or shallow spots where we thought they would be. One of the highlights of the hike was making our way up to the hab overview, which offered some amazing Martian views overlooking the hab.

Our second EVA was shortened due to strong winds that increased as the day went one. It was turned into a short relocation EVA where we dropped off our tiny friend “Paul Watney” back onto the vast plains of Mars.

The evening consisted of some more cooking for the crew, baking peanut butter cookies, and watching “Apollo 13.”

Look Ahead Plan: Finish cleaning Water Tank and final skirt HAB clean up. Exploration continues to the south and spreading North.

Anomalies in Work: None

Weather: Chilly early on, warmer later in the day, mostly cloudy all day.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report

Support Requested: None

Maintenance request: See Operations Report

Sol Summary – April 15th

Crew 192 Sol Summary Report 15Apr2018

Sol 7

Summary Title: It’s Not Possible, It’s Necessary

Author: Victoria Varone

Mission Status: Today was a half-stand down day for the crew.

Sol Activity Summary: One medical emergency simulation EVA in the afternoon, which taught the crew the basics of how to handle a medical emergency when out in the field on Mars. Crew doctor Ashok Narayanamoorthi shared the rules of what to do and what not to do, the most important being to remain calm.

The afternoon consisted of a debrief, some knot-tying lessons with the Commander, a Martian Matinee (we watched "Interstellar"), and some more fun cooking for dinner.

Look Ahead Plan: We’re aiming for some solid EVAs this week, exploring the stretches of the Martian terrain.

Anomalies in Work: None

Weather: Warm, partly sunny early in the day, cloudy in the afternoon.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report

Support Requested: None

Maintenance request: Almost out of bleach, could use a refill.

Sol Summary – April 14th

Crew 192 Sol Summary Report 14Apr2018

Sol 6

Summary Title: Here Comes the Sun

Author: Victoria Varone

Mission Status: The crew is now halfway through their mission! We were very excited to get back to EVAs after the last 2-3 days of strong winds and cold temperatures.

Sol Activity Summary: Today consisted of two really great EVAs for Crew 192. The weather started off cool and sunny and gradually became warmer, remaining clear and beautiful all day.

At the end of the second EVA, we started work on a hab repair request to remove the skirting from all the way around the hab. Some of it has been removed and the rest will be worked on during our second week in sim.

We ended up the day with an Italian meal of pasta and salad. Ciao bella!

Look Ahead Plan: Some stargazing at night (we’ll be using some red astronomy lights for night vision), a half-day’s break tomorrow plus a medical simulation EVA, some more planning for hab repairs and EVAs for next week.

Anomalies in Work: None

Weather: Warm and sunny.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Report, EVA Request Journalist Report, Photo Report

Support Requested: See Operations Report

Maintenance request: See Operations Report

Sol Summary – April 13th

Crew 192 Sol Summary Report 13Apr2018

Sol 5

Summary Title: The Winds of Mars

Author: Victoria Varone

Mission Status: Our Mars mission has been slightly inhibited by approximately two and a half days of strong winds (paging Mark Watney), but like most scientists, the crew is keeping themselves busy and happy finding other things to do (repairs, cleaning, organization, cooking) until the weather calms down.

Sol Activity Summary: With strong winds blowing again today, the crew attempted to gather a weather report before heading out on their first EVA of the day. Unfortunately, a sluggish connection prevented us from obtaining this prior to the EVA (Martian technology, right?). We left a little later than planned but were instructed to call off the outing, so we headed back inside to make a new plan for the morning and afternoon.

The winds started to pick up again a bit, so the crew used the extra time inside to conduct a day of cleaning and organizing the hab. This turned out to be a worthwhile task as we removed layers of dust from many areas and long-unused objects and organized some of the boxes and bins in the lower deck. After this, the crew enjoyed a delicious brunch of nachos crafted by their own Richard Blakeman.

We went ahead with our afternoon EVA as scheduled as the winds were slower than they were yesterday and our two crew members remained perfectly safe during their outing. Upon their return, we had a nice dinner, checked out the GreenHab together, looked through some of our many amazing mission photos so far, and planned for weekend EVAs and activities.

Look Ahead Plan: The forecast says that the winds will die down tonight and the weather will be beautiful over the weekend. We’ll use that time to make up for the EVAs and other activities that were prohibited over the last couple of days due to the strong winds.

Anomalies in Work: None

Weather: Chilly and windy. Winds blowing at approximately 23 mph with gusts 30 mph.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, EVA Report Journalist Report, Photo Report

Support Requested: See Operations Report

Maintenance request: See Operations Report

Sol Summary Report – 12 April 2018

Sol 4

Summary Title: Whatever It Takes

Author: Victoria Varone

Mission Status:

Sol Activity Summary: It was quite a busy day for Crew 192 as we spent almost all of our time literally “battening down the hatches” as severe winds caused the HAB some physical wear that needed repair and periodic observation.

It began at ~0220 when the wind peeked. We had to tie down the center wooden cap for the structure because it kept slamming against the ceiling and being pulled up by the wind. It is now secure.

We had an approved short morning scouting EVA …The commander recalled the team when the front airlock door vibrated and broke free from its hasp and the door swung open so hard that it broke the banister and damaged the habitat. The commander decided to have the team emergent repair the HAB by reattaching the hasp and screwing the banister back in place. The external repairs to the HAB must be made at a later date because they were not emergent and beyond the ability of the crew.

We completed another engineering EVA (ALL HANDS EFFORT) to clean the trailer water tank in the RAM so that our next crew can maintain a large and safe supply of water. This involved stringing 550 cord to make a safety line which allowed all personnel to get to the HAB without issue in the mounting wind.

Look Ahead Plan: The strong winds are forecast to continue into tomorrow and should slow down after that. Once it does, we will resume our exploration EVAs, continually exploring new sites and familiarizing ourselves further with ones we’ve already seen.

Anomalies in Work: None

Weather: Chilly and extremely windy. Winds blowing at approximately 30 mph with gusts 46 mph.

Crew Physical Status: All crew are physically healthy and safe.

Sol Summary – April 11th

Crew 192 Sol Summary Report

Sol 3

Summary Title: Gone With the Wind

Author: Victoria Varone

Mission Status: Today was the first really hot day of our mission, a surprise in between the last few days of cool weather and wind. We continued to explore further out into the Martian terrain, exercising extra caution this time to remain hydrated and not push ourselves past our limits in hazardous conditions. The search for our missing science experiments goes on and we simultaneously improvise new science experiments to execute in their place.

Sol Activity Summary: Our day started early with our first EVA, where crew members Joe Dituri and Andreaa Radulescu went to explore another site to do a bit of hiking and geological investigation. When they returned we all had lunch together, which was followed by a brainstorming session for our next few EVAs.

Then we relaxed and watched The Martian, which was surprisingly similar to our own mission (missing crew member, same number of remaining crew members, affinity for duck tape), especially since not long after, a crazy wind storm kicked up.

Upon hearing a lot of commotion outside, we decided to have crew members Ashok Narayanamoorthi and Richard Blakeman head out on an engineering EVA prior to our scheduled evening EVA to check all around the hab. They found that a railing near the main door was broken because the front air lock door swung open and smashed railing. They made a couple of adjustments to secure the remaining piece as well as the hab door. Following that, the wind became so strong and started to present a safety hazard to the crew so the leader of the away team (Richard Blakeman) terminated the EVA.

Look Ahead Plan: We will continue to extend our range of Martian exploration as far as we possibly can, but some forecast hazardous weather has prompted us to change up some of our EVA plans. With strong winds predicted for tomorrow and Friday, we may choose to instead remain closer to the hab and work on our water tank cleanup project in the RAM and resume exploration over the weekend. Safety ALWAYS comes first.

Anomalies in Work: None

Weather: Hot and windy. The temperature reached approximately 83 degrees today and the wind storm was severe.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, EVA Report Journalist Report, Commander’s Report, Photo Report, Science Report

Support Requested: See Commander’s Report.

Sol Summary – April 10th

Crew 192 Sol 2 Summary Report

Sol 2

Summary Title: A Beautiful Day for EVAs

Author: Victoria Varone

Mission Status: It was a beautiful day for EVAs over here for Crew 192 on Mars. We continue to clean and organize things around the hab and have started getting more comfortable with EVAs, which will allow us to continue the search for fascinating natural wonders as well as our missing science experiments and crew member. The mission is going well.

Sol Activity Summary: The day started with a morning EVA, where three of our crew checked out the Marble Ritual site and surrounding area. They discovered some interesting geological structures and were particularly fascinated by a view of snow-capped mountains from a dusty, sandy red desert. After exploring, the crew returned to the hab, examined one of the hab’s water tanks for future cleaning, and headed inside for lunch.

After a great afternoon session of debriefing, brainstorming, and improvising for additional science experiments, the remaining crew members headed out for another EVA, coming back with excited descriptions of their hike in Kissing Camels Ridge. They also had some insight into the length of time of EVAs, noting that the amount of hiking they did in their 2.5 hour EVA expended just enough of their energy without causing serious dehydration. They’ve advise the rest of us to limit high-impact EVAs to about two and a half hours for safety.

We’re now resting and planning for tomorrow’s activities and may end the night with a space-based movie.

Look Ahead Plan: The goal is continue stretching our EVAs a little further as we keep searching for our missing science experiments and crew member. Based on tonight’s hiking EVA, we also plan to pay special attention to crew health on EVAs, making sure everyone is healthy and comfortable and no one becomes dehydrated and ill. A little attention is also being put towards possible activities for our Yuri’s Night celebration.

Anomalies in Work: None

Weather: The weather was beautiful again today, the skies started out clear and the weather cool, which worked well for the first EVA. It became overcast from the afternoon on, which allowed our second EVA crew members to explore the sites with different skies overhead.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: None

Sol Summary – April 9th

Crew 192 Sol 1 Summary Report

Sol 1

Summary Title: PHEnOM Blue Crew Has Landed

Author: Victoria Varone

Mission Status: Ground Control to Major Tom, PHEnOM Blue Crew 192 has landed at the Mars Desert Research Station. Unfortunately we’ve “lost” a crew-member during reentry to Mars. Our science adviser was accidentally jettisoned along with many of our scientific experiments. We have renamed our habitat the Bowie Base in honor of the inspiration from the “star Man” himself David Bowie. Our team’s mission is to overcome this adversity and persevere in the name of science. We have good intelligence that some of the projects may have survived reentry so we will be performing EVAs to scout possible locations where the science may have re-entered. We intend to perform pilot studies and other scientific information gathering techniques as well.

Sol Activity Summary: We started our day by making sure we have an adequate water supply for our mission. Following that, we scouted the area around our habitat and took some photos for documentation. We also checked on our remaining science experiments like the locking mechanism of our observatory, which eventually opened after several unsuccessful attempts.

Next, we tested our ATVs and rovers to make sure they’re working properly for our EVAs and scouting missions. It was interesting getting to use them on the dusty, sandy Martian landscape.

After our test, we headed back inside the hab for our first crew dinner during our mission, and after that, two crew members conducted our first short-range EVA to test our spacesuits and communications devices. Upon their return, our GreenHab Officer went out to water the plants and our XO conducted a short engineering EVA around the hab for some repairs.

Look Ahead Plan: As we continue to search for our lost science experiments, we’ll stretch our EVAs further and further out into the Martian landscape each day. We’re also experimenting with various types of dehydrated “space food” to see what brings us the best combination of nutrition, convenience, and satisfaction. Some of the dehydrated items have proven to be more difficult to cook with than expected.

Anomalies in Work: None

Weather: The weather was beautiful all day today, clear skies with very few clouds. We’re getting a good feel for the desert climate, it’s dryness, and its variation in temperature. It seems to be somewhat chilly in the morning, warmer by midday (even a bit hot), and cools down nicely at night, sometimes even becoming very cold during the night itself. The dry air is also a change for some of the crew.

Crew Physical Status: All crew are physically healthy and safe.

EVA: Today two crewmembers conducted one short-range EVA that lasted approximately one hour. See additional EVA report.

Reports to be filed: Sol Summary Report, HSO Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: We would very much appreciate it if mission control could send us forecast sunset times (actual, civil, and nautical) for each day of our mission, which will allow us some better planning for EVAs.